OpenStudy (anonymous):

Use basic identities to simplify the expression tan/cot please help im not sure what identity to use

OpenStudy (chihiroasleaf):

cot = 1/tan

OpenStudy (anonymous):

so tan^2

OpenStudy (anonymous):

cot=cos/sine

OpenStudy (anonymous):

tan/cot=sin/cos*sin/c0s

OpenStudy (anonymous):

so i would be ending up with sin

OpenStudy (chihiroasleaf):

yes..., tan^2 ... :) what @ali110 write is also correct since we can write sin/cos*sin/c0s = sin^2/cos^2 = (sin/cos)^2 = tan^2

OpenStudy (anonymous):

sin^2/cos^2=(sin/cos)=tan^2

OpenStudy (anonymous):

ohhh ok yea @ali110 confused me alil

OpenStudy (anonymous):

but thanks alot guys <3

OpenStudy (anonymous):

i was writing it also but u interrupt!!! u fast girls:)

OpenStudy (anonymous):

hahaha thanks hun !!!

OpenStudy (chihiroasleaf):

you're welcome... :)

OpenStudy (anonymous):

@chihiroasleaf is so good in maths:)